OpenStudy (anonymous):

4/x=-20/15 solve for x in the proportion

hero (hero):

Hint: Cross Mulitply

OpenStudy (anonymous):

cross multiply

OpenStudy (anonymous):

ok

hero (hero):

@shivraj , why don't you post the result you get after cross multiplication.

OpenStudy (anonymous):

Here we can do it by another simple method..

hero (hero):

You mean a method other than cross multiplication?

OpenStudy (anonymous):

Yes..

OpenStudy (anonymous):

Let @shivraj solve first and at last I will tell the method..

hero (hero):

Okay, so @waterineyes will probably simplify the right fraction before cross multiplying. But that's nothing new.

OpenStudy (anonymous):

I will not use cross multiplication anywhere in my solution @Hero

hero (hero):

So how will your method be "simpler"?

OpenStudy (anonymous):

Can i show you??

hero (hero):

I can see that the fraction will be reduced to 4x = -4/3

hero (hero):

But now you still have to cross multiply. Doing anything else will just produce more steps.

OpenStudy (anonymous):

I won't i said @Hero

hero (hero):

err, I mean the fraction will be reduced to 4/x = -4/3 typo above

OpenStudy (anonymous):

Yes: \[\frac{4}{x} = \frac{-4}{3} \implies \frac{4}{x} = \frac{4}{-3} \implies x = -3\]

hero (hero):

Oh, I see what you did. Yeah, in that case, I guess you don't need to cross multiply. Nice trick.

OpenStudy (anonymous):

That is what I am trying to say..

hero (hero):

That's why the real first step is to actually simplify whenever possible

OpenStudy (anonymous):

Yes...

hero (hero):

Usually when I see a proportion, my first impulse is to cross multiply

OpenStudy (anonymous):

Ha ha ha..

OpenStudy (anonymous):

@waterineyes : Unknowingly you are cross multiplying the terms with 1. :) Still it's a good and simpler method. Well done.

OpenStudy (anonymous):

See the way to look this problem is : If; \[\frac{x}{y} = \frac{x}{z}\] then y must be equal to z that is why they become equal.. That is the trick here.. @Champs

OpenStudy (anonymous):

You are of course saying it mathematically but I am saying it logically...

hero (hero):

What do you call that? The Denominator Equivalency Property?

OpenStudy (anonymous):

Ha ha ha.. You can call that what you like..

hero (hero):

Or perhaps the Denominator Equality Property

OpenStudy (anonymous):

I like it. :)

OpenStudy (compassionate):

We see in this Problem that when a single fraction is equal to a single fraction, then the equation can be cleared by "cross-multiplying." If a c ---- = ---- b d , then ad = bc. -------------------------------------------------------- PROBLEM BELOW -------------------------------------------------------- x − 3 x − 5 ------ = ----- <=== SOLVE 3 2 Here is the cleared equation and its solution: 2(x − 3) = 3(x − 5) 2x − 6 = 3x − 15 2x − 3x = − 15 + 6 −x = −9 x = 9